# Deep Submillimeter and Radio Observations in the SSA22 Field. IV. Spectral Energy Distributions, Star Formation Histories, and the Infrared-Radio Correlation of the 850 -selected SMGs

### v1

@@ V.1 The IR-radio correlation factor
Blue inverted triangles and green squares in Figure 13 denote sources with an AGN fraction of at least 10% and those detected in X-rays, respectively. The AGN-tagged sources are predominantly outliers, exhibiting excess radio emission (Figure 12) and consequently lower values. These sources fall below or near the radio-excess AGNs threshold (yellow dashed line in Figure 13, 1.55; Algera et al., 2020), while several unmarked sources close to this criterion may also be potential AGN candidates. Interestingly, the sources exhibiting the lowest values are located around , potentially associated with the SSA22 protocluster (if consider photometric redshift uncertainties ). This suggests that dense environments may enhance galaxy interactions and/or leverage the abundant gas reservoirs within clusters promote galaxy evolution, thereby boosting AGN activity (Lehmer et al., 2009; Umehata et al., 2015). Among the AGN-flagged sources, only two have X-ray detections. Most X-ray-detected sources instead display near- or above-average , exhibiting elevated infrared luminosities. This implies that X-ray emission in DSFGs predominantly originates from vigorous star formation (e.g., X-ray binaries) rather than AGN activity.
Compared to local galaxy samples ( 2.64; Yun et al., 2001; Bell, 2003), our SMGs show systematically lower values by 0.27 dex, implying either reduced infrared luminosities or enhanced radio emission by a factor of 2. Nevertheless, the bulk of our SMG population remains within 2 of the local distribution (shaded region in left panel of Figure 13). These results align well with previous studies of infrared/radio-selected galaxies (Murphy et al., 2009; Ivison et al., 2009, 2010; Sargent et al., 2010a; Magnelli et al., 2015; Delhaize et al., 2017) and SMG samples (Murphy et al., 2009; Michałowski et al., 2010), which report median values in the range 2.34-2.41 and redshift evolution slopes between -0.10 and -0.26. Our findings show particularly close agreement with the large infrared/radio-selected sample of Magnelli et al. (2015); An et al. (2021), sharing similar distributions and evolutionary trends (slope -0.10 to -0.12) at 2.7-3 significance (Magnelli et al., 2015).
@@
We re-analyzed a sample from the SSA22 deep field, excluding sources likely affected by radio-loud AGN (i.e., those with fracAGN 0.1, mostly outliers) and high-redshift sources ( 4) where detection completeness may be an issue. For this “cleaned sample”, the median is 2.40 0.04 . A fit to this sample yields a redshift evolution slope = -0.07 0.03 (where (1+z)γ), with a significance of only 2.2. This indicates that, after removing AGN-dominated outliers, the overall evolutionary trend becomes flatter and less significant.
Our results are also consistent with those for SMG samples from Murphy et al. (2009); Michałowski et al. (2010). As illustrated in right panel of Figure 13, the fitting curve of total jointly submillimeter/radio selected sample (the black dotted line) exhibit an excellent agreement with the medain distribution reported by Michałowski et al. (2010) (green squares).
Michałowski et al. (2010) noted a significant difference in the median between sources at 1.5 and 1.5, despite claiming no strong evolution in SMGs overall. Motivated by this, we derived median values of this cleaned sample in each redshift bin (shown as red circles in right panel of Figure 13). For sources at 1.5, the median is 2.55 0.07, close to the local value (Bell, 2003). For the intermediate redshift bin 1.5-4, the median is 2.36 0.04 values are in close agreement with Michałowski et al. (2010) (see their Figure 6). Fitting the 1.5 and 1.5-4 subsamples separately yields slopes of 0.06 0.08 and -0.02 0.07, respectively, suggesting no clear evolution within either bin.
@@
We compare with both radio and infrared luminosities (Figure 14). Our results show that the trend of with 1.4 GHz radio luminosity () in SMGs is in good agreement with previous studies (Ivison et al., 2010; Sargent et al., 2010a), and this trend is also observed in nearby galaxies selected in the optical, infrared, and radio bands (Morić et al., 2010). A slight decline in is seen at [W Hz-1], which becomes more pronounced at higher luminosities and drops sharply for sources exceeding [W Hz-1]. Most sources above this radio luminosity threshold are flagged as AGN, suggesting that such high radio luminosities are likely dominated by powerful AGNs, which typically exhibit low values. In the highest infrared luminosity bin ( [W]), shows a tendency to increase with , consistent with Sargent et al. (2010a). This may indicate the presence of additional infrared emission mechanisms in galaxies with the most extreme star formation activity, or a relatively slower increase in radio emission compared to infrared. At [W], appears slightly higher than at intermediate luminosities, these sources are predominantly at , potentially reflecting the additional contribution from cirrus emission in low-redshift galaxies (Yun et al., 2001; Michałowski et al., 2010). However, this deviation is not statistically significant, and we conclude that is largely independent of infrared luminosity.
@@ V.1.1 Does the Infrared-Radio Correlation Evolve?
The most prominent factor may be AGN activity. Studies indicate a substantial AGN fraction within SMG samples (Alexander et al., 2005; Laird et al., 2010; Johnson et al., 2013). Delhaize et al. (2017) find that AGNs exhibit significantly lower values than star-forming galaxies (SFGs), with more luminous AGNs showing even lower and steeper evolutionary trends. As described above, AGN-tagged SMGs typically show excess radio emission and correspondingly lower . Moreover, the majority of SSA22 SMGs have 1.4 GHz luminosities exceeding [W Hz-1], a regime where the radio luminosity function is increasingly dominated by AGNs (Ivison et al., 2010). However, does AGN activity always lead to lower ? Murphy et al. (2009); Michałowski et al. (2010); Morić et al. (2010); Sargent et al. (2010a) note that most AGNs actually follow the same IRRC as star-forming galaxies, with values similar to those in the local universe. This implies that is ineffective at distinguishing obscured or radio-quiet AGNs (Murphy et al., 2009).
@@
Physical differences between high-redshift SMGs and local galaxies may lead to different values. High-redshift SMGs are not direct analogs of local ULIRGs (Menéndez-Delmestre et al., 2009). SMGs encompass a diverse population with a wide range of physical properties and a large intrinsic scatter in . Moreover, SMGs host extreme environments and physical processes. For example, Murphy et al. (2009) describe gas “bridges” forming between interacting SMG pairs (interactions and mergers are common in SMGs), which contain significant cosmic-ray electrons and magnetic fields, producing bright radio continuum emission. This radio emission can account for up to half of the system’s total radio output, even before substantial star formation is triggered in the bridge. Such systems exhibit excess radio emission, leading to low values. Additionally, extreme starburst systems in SMGs may possess mechanisms that enhance synchrotron efficiency. Michałowski et al. (2010) suggest that SMGs may be “puffy starbursts” (vertically and radially extended, with scale heights of 1 kpc), where reduced bremsstrahlung and ionization losses allow cosmic-ray electrons to produce stronger radio emission.
@@
Sample selection biases are also a critical factor that can result in an evolution trend of . Detection limitation mean that radio-selected samples miss radio-faint sources (Ivison et al., 2010), implying that such samples may preferentially include sources with stronger radio emission, leading to systematically lower values (Michałowski et al., 2010). Similarly, high-redshift infrared-selected samples are biased toward detecting the most luminous infrared galaxies (e.g., ULIRGs) (Sargent et al., 2010a), which may host more extreme star formation or AGN activity. As previously noted, Sargent et al. (2010a) demonstrate significant differences in between samples selected solely by radio, solely by infrared, and by combined detection, with these discrepancies growing at higher redshifts. The results of Casey et al. (2012) support this interpretation. Differences in may also stem from unaccounted mass selection biases. Delvecchio et al. (2021) argue that radio emission exhibits a stronger dependence on stellar mass than infrared emission, with more massive galaxies showing lower . Since SMGs are inherently biased toward massive systems, this could naturally explain differences in compared to local samples. The physical basis for a mass dependence could be that more massive galaxies confine cosmic-ray electrons more effectively, leading to higher non-thermal radio emission relative to their infrared output compared to lower-mass galaxies (Bell, 2003). It may also be related to star formation rate surface density (Delvecchio et al., 2021).
@@ VI Summary
Our SMG sample contributes, on average, 21% to the cosmic SFRD (Figure 9). This fraction peaks at 50-60% in the redshift range 2.5-3.5, but declines sharply at higher redshifts ( 4-6) due to changes in the star formation mode (Zavala et al., 2021). At lower redshifts, the contribution decreases further due to the “downsizing” effect, declining baryonic conversion efficiency, and wavelength selection biases. We also find that SMGs in overdense regions of the field make a significant contribution to the cosmic infrared background. The sample contributes 28% to the cosmic SMD, with a redshift evolution trend consistent with its SFRD contribution (Figure 10). The number density of SMGs evolves rapidly from the early universe to 1-3, increasing from 3-4 cMpc-3 to 4-6 cMpc-3 (Figure 11). At lower redshifts ( 1-1.5), 850 surveys miss a population of submillimeter-faint galaxies with warmer dust temperatures due to gas exhaustion, downsizing, and bandpass selection effects (Section IV.4).
@@ acknowledgments
Y.A. acknowledges the support from the National Key R&D Program of China (2023YFA1608204), the National Natural Science Foundation of China (NSFC grant 12173089), and the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No.XDB0800301).

### v2

@@ V IR-Radio Correlation
We must also caution about potential misidentifications in the sample, which are discussed in detail in Paper III. Sources with redshifts below may arise from spurious cross-matching. Moreover, since our sample identification relies heavily on radio counterparts, this selection could introduce additional systematic effects in the present analysis (Algera et al., 2020a). For instance, it is possible that some radio-bright, mid-infrared luminous but submillimeter-faint warm galaxies have been inadvertently included in our sample (Chapman et al., 2004). Additionally, it should be noted that we assumed a fixed radio spectral index of when converting 3 GHz flux densities to 1.4 GHz. While studies of large-sample SFGs suggest a median spectral index around with no strong redshift evolution (An et al., 2021), adopting a single value may introduce systematic errors. For instance, spectral aging could lead to an underestimation of high-frequency flux densities when extrapolating from lower frequencies (Thomson et al., 2019).
@@ V.1 The IR-Radio Correlation Factor
Blue inverted triangles and green squares in Figure 13 denote sources with an AGN fraction of at least 10% and those detected in X-rays, respectively. The AGN-tagged sources are predominantly outliers, exhibiting excess radio emission (Figure 12) and consequently lower values. These sources fall below or near the radio-excess AGNs threshold (yellow dashed line in Figure 13, 1.55; Algera et al., 2020b), while several unmarked sources close to this criterion may also be potential AGN candidates. Interestingly, the sources exhibiting the lowest values are located around , potentially associated with the SSA22 protocluster (if consider photometric redshift uncertainties ). This suggests that dense environments may enhance galaxy interactions and/or leverage the abundant gas reservoirs within clusters promote galaxy evolution, thereby boosting AGN activity (Lehmer et al., 2009; Umehata et al., 2015). Among the AGN-flagged sources, only two have X-ray detections. Most X-ray-detected sources instead display near- or above-average , exhibiting elevated infrared luminosities. This implies that X-ray emission in DSFGs predominantly originates from vigorous star formation (e.g., X-ray binaries) rather than AGN activity.
Compared to local galaxy samples ( 2.64; Yun et al., 2001; Bell, 2003), our SMGs show systematically lower values by 0.27 dex, implying either reduced infrared luminosities or enhanced radio emission by a factor of 2. Nevertheless, the bulk of our SMG population remains within 2 of the local distribution (shaded region in left panel of Figure 13). These results align well with previous studies of infrared/radio-selected galaxies (Murphy et al., 2009; Ivison et al., 2009, 2010; Sargent et al., 2010a; Magnelli et al., 2015; Delhaize et al., 2017) and SMG samples (Murphy et al., 2009), which report median values in the range 2.34-2.41 and redshift evolution slopes between -0.10 and -0.26. Our findings show particularly close agreement with the large infrared/radio-selected sample of Magnelli et al. (2015); An et al. (2021), sharing similar distributions and evolutionary trends (slope -0.10 to -0.12) at 2.7-3 significance (Magnelli et al., 2015).
@@
We re-analyzed a sample from the SSA22 deep field, excluding sources likely affected by radio-loud AGN (i.e., those with AGN fraction 0.1, mostly outliers) and high-redshift sources ( 4) where detection completeness may be an issue. For this “cleaned sample”, the median is 2.40 0.04 . A fit to this sample yields a redshift evolution slope = -0.07 0.03 (where (1+z)γ), with a significance of only 2.2 (red solid line in right panel of Figure 13). This indicates that, after removing AGN-dominated outliers, the overall evolutionary trend becomes flatter and less significant.
Our results are also consistent with those for SMG samples from Murphy et al. (2009). Michałowski et al. (2010) noted a significant difference in the median between sources at 1.5 and 1.5, despite claiming no strong evolution in SMGs overall. Motivated by this, we derived median values of this cleaned sample in each redshift bin (shown as red circles in right panel of Figure 13). For sources at 1.5, the median is 2.55 0.07, close to the local value (Bell, 2003). For the intermediate redshift bin 1.5-4, the median is 2.36 0.04 values are in close agreement with Michałowski et al. (2010) (see their Figure 6). Fitting the 1.5 and 1.5-4 subsamples separately yields slopes of 0.06 0.08 and -0.02 0.07, respectively (blue dashed lines in right panel of Figure 13), suggesting no clear evolution within either bin.
@@
For radio-undetected sources via stacking analysis, Thomson et al. (2014) derived a median of for the ALESS sample. After further restricting the sample to and removing radio-excess SMGs, the median remains at 2.33, which is remarkably close to our results. In contrast, the large-sample study of SMGs from the AS2UDS survey reported a lower median of , which shows no evolution with redshift or other physical properties (Algera et al., 2020a). The authors argue that SMGs exhibit a different IRRC compared to local galaxies, likely due to a combination of strong magnetic fields, high interstellar medium densities, and additional radio emission from secondary cosmic rays. Local ULIRGs share comparable magnetic field strengths with SMGs but reside in even denser environments. Thomson et al. (2019) pointed out that spectral aging, manifesting as convex spectral behavior, is present in some SMGs. This effect may result from young central starbursts and suppressed radio emission due to weaker magnetic fields, and could serve as evidence for evolution. Indeed, Thomson et al. (2014) noted that co-evolves with the radio spectral index , offering a potential tracer of the evolutionary stage of high-redshift starburst galaxies. A lower obscured fraction of star formation at early cosmic times may also contribute to the evolution of . As highlighted by Shen et al. (2022), while radio emission can respond on short timescales, dust buildup and stellar population evolution operate over significantly longer timescales.
We compare with both radio and infrared luminosities (Figure 14). Our results show that the trend of with 1.4 GHz radio luminosity () in SMGs is in good agreement with previous studies (Ivison et al., 2010; Sargent et al., 2010a), and this trend is also observed in nearby galaxies selected in the optical, infrared, and radio bands (Morić et al., 2010). A slight decline in is seen at [W Hz-1], which becomes more pronounced at higher luminosities and drops sharply for sources exceeding [W Hz-1]. Most sources above this radio luminosity threshold are flagged as AGN, suggesting that such high radio luminosities are likely dominated by powerful AGNs, which typically exhibit low values. In the highest infrared luminosity bin ( [W]), shows a tendency to increase with , consistent with Sargent et al. (2010a). This may indicate the presence of additional infrared emission mechanisms in galaxies with the most extreme star formation activity, or a relatively slower increase in radio emission compared to infrared. At [W], appears slightly higher than at intermediate luminosities, these sources are predominantly at , potentially reflecting the additional contribution from cirrus emission in low-redshift galaxies (Yun et al., 2001). However, this deviation is not statistically significant, and we conclude that is largely independent of infrared luminosity.
@@ V.1.1 Does the Infrared-Radio Correlation Evolve?
The most prominent factor may be AGN activity. Studies indicate a substantial AGN fraction within SMG samples (Alexander et al., 2005; Laird et al., 2010; Johnson et al., 2013). Delhaize et al. (2017) find that AGNs exhibit significantly lower values than star-forming galaxies (SFGs), with more luminous AGNs showing even lower and steeper evolutionary trends. By performing morphological classification of optically selected galaxies at relatively low redshifts (), Molnár et al. (2017) find that spheroid-dominated systems exhibit significant evolution, and they are more likely to be contaminated by AGN. As described above, AGN-tagged SMGs typically show excess radio emission and correspondingly lower . Moreover, the majority of SSA22 SMGs have 1.4 GHz luminosities exceeding [W Hz-1], a regime where the radio luminosity function is increasingly dominated by AGNs (Ivison et al., 2010). However, does AGN activity always lead to lower ? Murphy et al. (2009); Morić et al. (2010); Sargent et al. (2010a) note that most AGNs actually follow the same IRRC as star-forming galaxies, with values similar to those in the local universe. This implies that is ineffective at distinguishing obscured or radio-quiet AGNs (Murphy et al., 2009).
@@
Physical differences between high-redshift SMGs and local galaxies may lead to different values. High-redshift SMGs are not direct analogs of local ULIRGs (Menéndez-Delmestre et al., 2009). SMGs encompass a diverse population with a wide range of physical properties and a large intrinsic scatter in . Moreover, SMGs host extreme environments and physical processes. For example, Murphy et al. (2009) describe gas “bridges” forming between interacting SMG pairs (interactions and mergers are common in SMGs), which contain significant cosmic-ray electrons and magnetic fields, producing bright radio continuum emission. This radio emission can account for up to half of the system’s total radio output, even before substantial star formation is triggered in the bridge. Such systems exhibit excess radio emission, leading to low values. Additionally, extreme starburst systems in SMGs may possess mechanisms that enhance synchrotron efficiency. Michałowski et al. (2010) suggest that SMGs may be “puffy starbursts” (vertically and radially extended, with scale heights of 1 kpc), where reduced bremsstrahlung and ionization losses allow cosmic-ray electrons to produce stronger radio emission. Thomson et al. (2019) suggested that a young central starburst combined with a weak magnetic field can suppress radio emission, leading to spectral aging. Meanwhile, Thomson et al. (2014) proposed that the infrared-to-radio ratio co-evolves with the radio spectral index alpha. Additionally, an initially low obscuration fraction during early star formation may also contribute to the observed evolution of . As citetshen2022 pointed out, radio emission can respond on short timescales, whereas dust buildup and stellar population maturation require significantly longer timescales.
@@
Sample selection biases are also a critical factor that can result in an evolution trend of . Detection limitation mean that radio-selected samples miss radio-faint sources (Ivison et al., 2010), implying that such samples may preferentially include sources with stronger radio emission, leading to systematically lower values. Similarly, high-redshift infrared-selected samples are biased toward detecting the most luminous infrared galaxies (e.g., ULIRGs) (Sargent et al., 2010a), which may host more extreme star formation or AGN activity. As previously noted, Sargent et al. (2010a) demonstrate significant differences in between samples selected solely by radio, solely by infrared, and by combined detection, with these discrepancies growing at higher redshifts. The results of Casey et al. (2012) support this interpretation. Differences in may also stem from unaccounted mass selection biases. Delvecchio et al. (2021) argue that radio emission exhibits a stronger dependence on stellar mass than infrared emission, with more massive galaxies showing lower . Since SMGs are inherently biased toward massive systems, this could naturally explain differences in compared to local samples. The physical basis for a mass dependence could be that more massive galaxies confine cosmic-ray electrons more effectively, leading to higher non-thermal radio emission relative to their infrared output compared to lower-mass galaxies (Bell, 2003). It may also be related to star formation rate surface density (Delvecchio et al., 2021).
@@ V.2 Submm/Radio Flux Ratio
However, we caution that potential misidentification issues may exist. First, sources below redshift 1 could arise from incorrect cross-identification (Paper III). Second, since our sample identification relies on radio-band counterparts, we may inadvertently include some galaxies that are faint in the submillimeter but bright in the radio and mid-infrared bands, such as warm, radio-loud systems (Chapman et al., 2004).
@@ VI Summary
Our SMG sample contributes, on average, 21% to the cosmic SFRD (Figure 9). This fraction peaks at 50-60% in the redshift range 2.5-3.5, but declines sharply at higher redshifts ( 4-6) due to changes in the star formation mode (Bourne et al., 2017; Dunlop et al., 2017; Bouwens et al., 2020, 2022; Zavala et al., 2021). At lower redshifts, the contribution decreases further due to the “downsizing” effect, declining baryonic conversion efficiency, and wavelength selection biases. We also find that SMGs in overdense regions of the field make a significant contribution to the cosmic infrared background. The sample contributes 28% to the cosmic SMD, with a redshift evolution trend consistent with its SFRD contribution (Figure 10). The number density of SMGs evolves rapidly from the early universe to 1-3, increasing from 3-4 cMpc-3 to 4-6 cMpc-3 (Figure 11). At lower redshifts ( 1-1.5), 850 surveys miss a population of submillimeter-faint galaxies with warmer dust temperatures due to gas exhaustion, downsizing, and bandpass selection effects (Section IV.4).
@@ acknowledgments
We sincerely thank Ian Smail for his constructive comments and insightful suggestions, which significantly improved the rigor and clarity of this manuscript.
Y.A. acknowledges the support from the National Key R&D Program of China (2023YFA1608204), the National Natural Science Foundation of China (NSFC grant 12173089), the China Manned Space Program with grant no. CMS-CSST-2025-A10, and the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No.XDB0800301).